OpenStudy (anonymous):

X to 2nd power+6x=-3 Find a, b, and c It will help you to put the equation in standard form. Solve equation using quadratic formula.

OpenStudy (anonymous):

x^2 +6x=-3 standard form ax^2+bx+c x^2+6x+3 so look at corresponding places a=1 b=6 c=3

OpenStudy (anonymous):

a=1, b=6, c=3. \[-6\pm \sqrt{6^{2}-4(1)(3)}/2(1)\]

OpenStudy (anonymous):

There's your set-up.

OpenStudy (anonymous):

Thanks for the answers, but I still don't know how to solve the equation.

OpenStudy (anonymous):

Um... why not?

OpenStudy (anonymous):

I get to -6 + sqrt of 12. and then I think sqrt of 12 is 3.464. then -6 + sqrt of 3.464. doesn't make sense

OpenStudy (anonymous):

When I do the formula I get -6 plus7minus sqrt(36-12) That what you got?

OpenStudy (anonymous):

"plus/minus" not plus7minus

OpenStudy (anonymous):

right and it's all divided by 2. so that makes -3 (+ or -) sqrt of 12. right?

OpenStudy (anonymous):

right and it's all divided by 2. so that makes -3 (+ or -) sqrt of 12. right?

OpenStudy (anonymous):

\[-3\pm \sqrt{12}\]

OpenStudy (anonymous):

Sorry, had to go out. Nearly, you can't just divide the 24 by 2 because the 24 is under a square root sign and your 2 isn't. You need to factor a 4 out of the square root (which becomes a 2 when you take it out, hope you can see this). Then u can cancel your 2 with that two, leaving -3 plus/minus sqrt 6.